OpenStudy (anonymous):

Prove if x>0, y>0 and x

OpenStudy (anonymous):

Consider variables x and y, where y is n larger than x. y = x+n Where n is any positive real number. x and y are real numbers greater than 1. This satisfies your conditions of x<y and x>0, y>0. Therefore: \[y^2 = (x+n)^2\] expanding the right side: \[y^2 = x^2 + nx + n^2\] Therefore, for the assertion that x^2 < y^2 is true, then the following must also be true: \[x^2 < x^2 + nx + n^2\] The range of nx and n^2 for any positive, real n and any x>1 is always positive, therefore it is true.

OpenStudy (shubhamsrg):

y^2 - x^2 = (x+y)(y-x) now (x+y)(y-x) > 0 always, since x>0 and y>0 also y>x which means y^2 - x^2 > 0 or y^2 > x^2 ..

OpenStudy (anonymous):

Same as @shubhamsrg Given: y>x y-x>0 (y+x)(y-x)>(y+x)*0 [ since x>0 ,y>0 (x+y)>0] (y^2-x^2)>0 y^2>x^2

OpenStudy (anonymous):

[ since x>0 ,y>0 (x+y)>0]* the inequality symbol doesnot changes

OpenStudy (anonymous):

thanks guys!